# Model Already Knows the Best Noise: Bayesian Active Noise Selection via Attention in Video Diffusion Model

### v1

@@ Abstract
The choice of initial noise significantly affects the quality and prompt alignment of video diffusion models, where different noise seeds for the same prompt can lead to drastically different generations. While recent methods rely on externally designed priors such as frequency filters or inter-frame smoothing, they often overlook internal model signals that indicate which noise seeds are inherently preferable. To address this, we propose ANSE (Active Noise Selection for Generation), a model-aware framework that selects high-quality noise seeds by quantifying attention-based uncertainty. At its core is BANSA (Bayesian Active Noise Selection via Attention), an acquisition function that measures entropy disagreement across multiple stochastic attention samples to estimate model confidence and consistency. For efficient inference-time deployment, we introduce a Bernoulli-masked approximation of BANSA that enables score estimation using a single diffusion step and a subset of attention layers. Experiments on CogVideoX-2B and 5B demonstrate that ANSE improves video quality and temporal coherence with only an 8% and 13% increase in inference time, respectively, providing a principled and generalizable approach to noise selection in video diffusion. See our project page: https://anse-project.github.io/anse-project/
@@
$\dagger$\$$\dagger$\$footnotetext: First and corresponding author
Figure 2: Conceptual comparison of noise initialization. (a) Prior methods (1; 2) iteratively refine noise using frequency domain priors through full diffusion sampling, incurring significant computational cost. (b) In contrast, our approach selects optimal noise seeds by estimating attention-based uncertainty at the first denoising step, enabling efficient and model-aware noise selection.
@@ 1 Introduction
Diffusion models have rapidly established themselves as a powerful class of generative models, demonstrating state-of-the-art performance across images and videos (3; 4; 5; 6; 7; 8; 9; 10; 11). In particular, Text-to-Video (T2V) diffusion models have received increasing attention for their ability to generate temporally coherent and visually rich video sequences. To achieve this, most T2V model architectures extend Text-to-Image (T2I) diffusion backbones by incorporating temporal modules or motion-aware attention layers (12; 13; 14; 15; 7; 16; 8). Furthermore, other works explore video generative structures, such as causal autoencoders or video autoencoder-based models, which aim to generate full video volumes rather than a sequence of independent frames (17; 9; 18; 19; 10; 11).
Beyond architectural design, another promising direction lies in improving noise initialization at inference time for T2I and T2V generation (20; 21; 22; 23). This aligns with the growing trend of inference-time scaling, observed not only in Large Language Models (24; 25) but also in diffusion-based generation systems (26). Due to the iterative nature of the diffusion process, the choice of initial noise profoundly influences video quality, temporal consistency, and prompt alignment (27; 1; 28; 2). As illustrated in Figure 1, the same prompt can lead to drastically different videos depending solely on the noise seed, motivating the need for intelligent noise selection.
Several recent approaches attempt to address this by designing external noise priors. For example, PYoCo (27) introduces inter-frame dependent noise patterns to improve coherence, though it requires extensive fine-tuning. FreeNoise (28) reschedules noise across time using a fusion-based strategy, while FreeInit (1) applies frequency-domain filtering to preserve low-frequency components. FreqPrior (2) extends this idea via Gaussian-shaped frequency priors and partial sampling. While effective, these methods rely on externally designed priors and require multiple full diffusion passes to evaluate candidate seeds. More importantly, they fail to leverage internal signals within the model that indicate which noise seeds are inherently preferable.
@@
Unlike BALD(29), which operates on classification logits, BANSA measures entropy in attention maps, arguably the most informative signals in generative diffusion. It compares the average entropy of individual maps to the entropy of their mean, capturing both uncertainty and disagreement across forward passes. A low BANSA score indicates that the model is more confident and certain in its attention, which empirically correlates with coherent video generation, as shown in Figure 1.
To make this approach suitable for inference-time, we approximate BANSA using Bernoulli-masked attention, which enables multiple stochastic attention samples from a single forward pass. We further reduce computation by limiting BANSA evaluation to early denoising steps and a subset of informative attention layers, selected via correlation analysis. Our contributions are threefold:
  • We present ANSE, the first active noise selection framework for video diffusion models, built on a principled Bayesian formulation of attention-based uncertainty.
  • We introduce BANSA, a novel acquisition function that quantifies attention consistency under stochastic perturbations, enabling model-aware noise selection without retraining or external noise priors.
  • Our method enhances both video quality and temporal consistency across various text-to-video architectures, with only a marginal increase in inference time of about 8% for CogVideoX-2B and 13% for CogVideoX-5B.
@@ 2 Preliminary
Diffusion models (30; 31) have achieved remarkable success across generative tasks. For T2V generation, directly operating in pixel space incurs high computational cost. To address this, video diffusion models (VDMs) typically adopt the latent diffusion model (LDM) framework, where the diffusion process is conducted in a compressed latent space.
A video autoencoder, composed of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, is trained to reconstruct the original input video $\mathbf{x}$ such that $\mathbf{x}=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. Denoting the latent code as $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x})$, the forward diffusion process adds noise over time:
@@
where $\bar{\alpha}_{t}$ is a pre-defined variance schedule. To learn the reverse process, a denoising network $\boldsymbol{\epsilon}_{\theta}$ is trained using the denoising score matching loss (32):
@@
where $\mathbf{c}$ denotes the conditioning text prompt. During sampling, the generation begins from Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and proceeds via a deterministic DDIM solver (33). The update at each step is computed as:
@@
where the denoised latent estimate $\hat{\mathbf{z}}_{0}(t):=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c},t)}{\sqrt{\bar{\alpha}_{t}}}$ is obtained using Tweedie’s formula (34; 35). This iterative process continues until $t=1$, yielding the final denoised latent $\mathbf{z}_{0}$, which is decoded into a video via $\mathcal{D}$.
@@
Active Learning improves model performance by selecting the most informative samples from an unlabeled pool in training phase. Acquisition functions are typically categorized into uncertainty-based (36; 29; 37; 38) and distribution-based (39; 40; 41; 42) approaches, with some relying on external modules such as auxiliary predictors (38; 43; 44). While active learning has been predominantly applied to image classification tasks, in this work, we focus on adapting uncertainty-based methods to text-to-video generation, without requiring additional models.
Predictive entropy is a common uncertainty measure, but it captures only aleatoric uncertainty and fails to account for parameter uncertainty. BALD addresses this by quantifying epistemic uncertainty via the mutual information between predictions $\mathbf{y}$ and model parameters $\theta$:
@@
where $\mathcal{H}[p]=-\sum_{y}p(y)\log p(y)$ is the Shannon entropy (45). The first term captures the entropy of the mean prediction, while the second term averages the entropy over stochastic forward passes. A high BALD score indicates confident but disagreeing predictions, revealing high epistemic uncertainty.
Since the posterior over $\theta$ is intractable, BALD is approximated using $K$ stochastic forward passes (e.g., Monte Carlo dropout):
@@ 3 Methods
We propose ANSE, a framework for selecting high-quality noise seeds in T2V diffusion models based on model uncertainty as shown in Figure 2. ANSE is built upon an acquisition function called BANSA, which extends uncertainty-based criteria from classification tasks to the attention space of generative diffusion models (Section 3.1). To enable efficient inference-time application, we approximate BANSA using Bernoulli-masked attention sampling (Section 3.2). Furthermore, to reduce computational redundancy, we identify a representative attention layer using correlation-based linear probing (Section 3.3). The overall pipeline is illustrated in Figure 3.
@@ 3.1 BANSA: Bayesian Active Noise Selection via Attention
We introduce BANSA, an acquisition function for selecting optimal noise seeds in T2V diffusion models. Unlike classification tasks with explicit predictive distributions, diffusion models lack such outputs. We instead estimate uncertainty in the attention space, where alignment between text and visual tokens naturally emerges during generation. Here, attention maps are treated as stochastic predictions conditioned on the noise seed $\mathbf{z}$, prompt $c$, and diffusion timestep $t$. BANSA measures disagreement and confidence across multiple attention samples, capturing attention-level uncertainty analogous to BALD, but tailored to the generative setting.
@@
$\text{BANSA}(\mathbf{z},c,t):=\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)})-\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right),$ (4)
where the row-wise entropy of an attention map $A$ is given by:
$\mathcal{H}(A):=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}-A_{ij}\log A_{ij}.$ (5)
This formulation captures both the sharpness (confidence) and the consistency (agreement) of attention behavior. BANSA can be applied to various attention types (e.g., cross-, self-, or temporal) and allows layer-wise interpretability.
Given a noise pool $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$, we select the optimal noise seed that minimizes the BANSA score:
@@
The proof is deferred to the Appendix. This condition implies that minimizing the BANSA score promotes attention behavior that is both confident and consistent under stochastic perturbations. Empirically, such attention patterns are associated with better prompt alignment, temporal coherence, and visual fidelity in generated videos. Therefore, BANSA serves as a principled criterion for model-aware noise selection in T2V.
@@ 3.2 Stochastic Approximation of BANSA via Bernoulli-Masked Attention
While BANSA provides a principled objective for noise selection, its computation requires $K$ independent forward passes per noise seed $\mathbf{z}$, which is computationally expensive. To mitigate this cost, we propose a stochastic approximation using Bernoulli-masked attention, enabling multiple attention samples from a single pass.
Instead of computing stochastic attention maps from $K$ separate forward passes such as equipping dropout, we inject stochasticity directly into the attention computation by applying binary masks to the attention scores. For each sample iteration $k=1,\cdots,K$, we generate a binary mask $m_{k}\in\{0,1\}^{N\times N}$ where each element is drawn i.i.d. from $\text{Bernoulli}(p)$. The masked attention map is computed as:
$\hat{A}^{(k)}(\mathbf{z},c,t):=\texttt{Softmax}\left(\left(\mathbf{Q}(\mathbf{z},c,t)\,\mathbf{K}(\mathbf{z},c,t)^{\top}\right)\odot m_{k}\right).$ (7)
where $\odot$ denotes element-wise multiplication. These masks simulate variability in attention patterns while keeping the input $(\mathbf{z},c,t)$ fixed. Using $K$ such samples, we define the approximate BANSA:
[…]
@@
5 Generate video: $\hat{v}=\texttt{SampleVideo}(\mathbf{z}^{*},c,t)$;
return $\hat{v}$
@@ 3.3 Layer Selection via Cumulative BANSA Correlation
BANSA can be computed at any attention layer, but attention behavior varies across depth. While using all layers provides a comprehensive uncertainty estimate, it is computationally heavy for deep T2V models. To address this, we propose a correlation-based truncation strategy that selects the smallest depth $d^{*}$ such that the averaged BANSA score over the first $d$ layers remains highly correlated with the full-layer score.

### v2

# arXiv:2505.17561v2[cs.CV] 01 Mar 2026

@@ Abstract
The choice of initial noise strongly affects quality and prompt alignment in video diffusion; different seeds for the same prompt can yield drastically different results. While recent methods use externally designed priors (e.g., frequency filtering or inter-frame smoothing), they often overlook internal model signals that indicate inherently preferable seeds. To address this, we propose ANSE (Active Noise Selection for Generation), a model-aware framework that selects high-quality seeds by quantifying attention-based uncertainty. At its core is BANSA (Bayesian Active Noise Selection via Attention), an acquisition function that measures entropy disagreement across multiple stochastic attention samples to estimate model confidence and consistency. For efficient inference-time deployment, we introduce a Bernoulli-masked approximation of BANSA that estimates scores from a single diffusion step and a subset of informative attention layers. Experiments across diverse text-to-video backbones demonstrate improved video quality and temporal coherence with marginal inference overhead, providing a principled and generalizable approach to noise selection in video diffusion. See our project page: https://anse-project.github.io/anse-project/.
@@
[…]
Figure 2: Conceptual comparison of noise initialization. (a) Prior methods (Wu et al., 2024; Yuan et al., 2025) refine noise via frequency priors and full diffusion sampling, leading to high cost. (b) Our method instead selects noise seeds by estimating attention-based uncertainty at the first denoising step, enabling efficient, model-aware selection.
@@ 1 Introduction
Diffusion models have rapidly established themselves as a powerful class of generative models, demonstrating state-of-the-art performance across images and videos (Rombach et al., 2022; Esser et al., 2024; Xie et al., 2024; Chen et al., 2024b; a; Wang et al., 2025; Yang et al., 2024; Zheng et al., 2024; Team, 2025). In particular, Text-to-Video (T2V) diffusion models have received increasing attention for their ability to generate temporally coherent and visually rich video sequences. To achieve this, most T2V model architectures extend Text-to-Image (T2I) diffusion backbones by incorporating temporal modules or motion-aware attention layers (Blattmann et al., 2023; He et al., 2022; Wang et al., 2023; Chen et al., 2023; 2024a; Guo et al., 2024b; Wang et al., 2025). Furthermore, other works explore video generative structures, such as causal autoencoders or video autoencoder-based models, which aim to generate full video volumes rather than a sequence of independent frames (Hong et al., 2022; Yang et al., 2024; HaCohen et al., 2024; Kong et al., 2024; Zheng et al., 2024; Team, 2025).
Beyond architectural design, another promising direction lies in improving noise initialization at inference time for T2I and T2V generation (Guo et al., 2024a; Eyring et al., 2024; Chen et al., 2024c; Xu et al., 2025). This aligns with the growing trend of inference-time scaling, observed not only in Large Language Models (Brown et al., 2024; Snell et al., 2024) but also in diffusion-based generation systems (Ma et al., 2025). Due to the iterative nature of the diffusion process, the choice of initial noise profoundly influences video quality, temporal consistency, and prompt alignment (Ge et al., 2023; Wu et al., 2024; Qiu et al., ; Yuan et al., 2025). As illustrated in Figure 1, the same prompt can lead to drastically different videos depending solely on the noise seed, motivating the need for intelligent noise selection.
Recent approaches tackle this by introducing external noise priors. PYoCo (Ge et al., 2023) enforces inter-frame dependent patterns for coherence but requires heavy fine-tuning. FreeNoise (Qiu et al., ) reschedules noise across time with a fusion strategy, FreeInit (Wu et al., 2024) preserves low-frequency components via frequency filtering, and FreqPrior (Yuan et al., 2025) extends this with Gaussian-shaped priors and partial sampling. While effective, these methods rely on external priors and repeated full diffusion passes, while ignoring internal model signals that identify inherently preferable seeds.
@@
Unlike BALD (Gal et al., 2017), which estimates uncertainty from classification logits, BANSA measures entropy in attention maps, arguably the most informative signals in diffusion. It compares the mean of per-pass entropies with the entropy of the mean map, capturing both uncertainty and cross-pass disagreement. A lower BANSA score indicates more confident, consistent attention and empirically correlates with more coherent video generation (Fig. 1). To make BANSA inference-friendly, we approximate it with Bernoulli-masked attention, yielding multiple stochastic attention samples from a single forward pass. We further cut cost by evaluating early denoising steps and only a subset of informative layers selected via correlation analysis. Our contributions are threefold:
  • We present ANSE, the first active noise selection framework for video diffusion, grounded in a Bayesian formulation of attention-based uncertainty.
  • We introduce BANSA, an acquisition function that measures attention consistency under stochastic perturbations, enabling model-aware noise selection without retraining.
  • Our method improves video quality and temporal consistency across diverse text-to-video backbones, with only marginal inference overhead.
@@ 2 Preliminary
Diffusion models (Ho et al., 2020; Song et al., 2021b) have achieved strong results across generative tasks. In T2V, pixel-space diffusion is costly, so most video diffusion models adopt latent diffusion and operate in a compressed latent space. A video autoencoder with encoder $\mathcal{E}$ and decoder $\mathcal{D}$ reconstructs $\mathbf{x}$ as $\mathbf{x}=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. Let $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x})$. The forward diffusion then progressively adds noise over time:
@@
where $\bar{\alpha}_{t}$ is a pre-defined variance schedule. To learn the reverse process, a denoising network $\boldsymbol{\epsilon}_{\theta}$ is trained using the denoising score matching loss (Vincent, 2011):
@@
where $\mathbf{c}$ is the conditioning text. Sampling starts from Gaussian noise $\mathbf{z}_{T}\!\sim\!\mathcal{N}(0,\mathbf{I})$ and uses a deterministic DDIM solver (Song et al., 2021a). Each step updates as:
@@
where the denoised latent estimate $\hat{\mathbf{z}}_{0}(t):=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c},t)}{\sqrt{\bar{\alpha}_{t}}}$ is obtained using Tweedie’s formula (Efron, 2011; Kim & Ye, 2021). This iterative process continues until $t=1$, yielding the final denoised latent $\mathbf{z}_{0}$, which is decoded into a video via $\mathcal{D}$.
@@
Active learning selects the most informative samples from an unlabeled pool to improve training. Acquisition functions are commonly uncertainty-based (Houlsby et al., 2011; Gal et al., 2017; Kirsch et al., 2019; Yoo & Kweon, 2019) or distribution-based (Mac Aodha et al., 2014; Yang et al., 2015; Sener & Savarese, 2017; Sinha et al., 2019), and some use external modules such as auxiliary predictors (Yoo & Kweon, 2019; Tran et al., 2019; Kim et al., 2021). While most prior work targets image classification, we adapt uncertainty-based acquisition to text-to-video generation without additional models.
Predictive entropy is widely used, but it reflects data noise and does not isolate parameter uncertainty. BALD instead measures epistemic uncertainty via the mutual information between predictions $\mathbf{y}$ and parameters $\theta$:
@@
where $\mathcal{H}[p]=-\sum_{y}p(y)\log p(y)$ is Shannon entropy (Shannon, 1948). The first term is the entropy of the mean prediction, and the second is the average entropy across stochastic forward passes. A high BALD score means predictions are confident yet disagree, indicating high epistemic uncertainty. Since the posterior over $\theta$ is intractable, BALD is approximated using $K$ stochastic forward passes (e.g., Monte Carlo dropout):
@@ 3 Methods
We propose ANSE, a framework for selecting high-quality noise seeds in T2V diffusion based on model uncertainty (Fig. 2). ANSE centers on BANSA, an acquisition function that transfers uncertainty criteria from classification to the attention space of generative diffusion models (Sec. 3.1). For efficient inference, we approximate BANSA via Bernoulli-masked attention sampling (Sec. 3.2). To avoid redundancy, we select a representative attention layer using correlation-based linear probing (Sec. 3.3). The full pipeline appears in Fig. 3.
@@ 3.1 BANSA: Bayesian Active Noise Selection via Attention
We introduce BANSA, an acquisition function for selecting optimal noise seeds in T2V diffusion. Unlike classifiers with explicit predictive distributions, diffusion models do not expose such outputs, so we estimate uncertainty in the attention space where text and visual tokens align during generation. We treat attention maps as stochastic predictions conditioned on the seed $\mathbf{z}$, prompt $\mathbf{c}$, and timestep $t$. BANSA measures both disagreement and confidence across multiple attention samples, providing a BALD-style uncertainty criterion tailored to the generative setting.
@@
$\text{BANSA}(\mathbf{z},c,t):=\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right)-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)}),$ (4)
where $\mathcal{H}(A):=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}-A_{ij}\log A_{ij}.$ This formulation captures both the sharpness (confidence) and the consistency (agreement) of attention behavior. BANSA can be applied to various attention types (e.g., cross-, self-, or temporal) and allows layer-wise interpretability.
Given a noise pool $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$, we select the optimal noise seed that minimizes score:
@@
The proof is deferred to the Appendix B. This result implies that minimizing the BANSA score encourages attention that is both confident and consistent under stochastic perturbations. Empirically, low BANSA correlates with better prompt alignment, temporal coherence, and visual fidelity. Thus, BANSA provides a principled, model-aware criterion for noise selection in T2V.
@@ 3.2 Stochastic Approximation of BANSA via Bernoulli-Masked Attention
While BANSA is principled for noise selection, computing it with $K$ independent passes per seed $\mathbf{z}$ is costly. We instead draw $K$ stochastic attention samples from a single pass via Bernoulli-masked attention. Rather than running $K$ passes, we inject stochasticity directly into the attention scores. For each $k\!=\!1,\dots,K$, sample a binary mask $m_{k}\in\{0,1\}^{N\times N}$ i.i.d. from $\mathrm{Bernoulli}(p)$ and compute masked attention map, $\hat{A}^{(k)}(\mathbf{z},c,t):={A}(\mathbf{z},c,t)\odot m_{k},$ $\odot$ denotes element-wise multiplication and rows are re-normalized. Using these $K$ such samples, we define the approximate BANSA:
${\text{BANSA-E}}(\mathbf{z},c,t):=\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}\hat{A}^{(k)}(\mathbf{z},c,t)\right)-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(\hat{A}^{(k)}(\mathbf{z},c,t)).$ (6)
By concavity of entropy, $\text{BANSA-E}\geq 0$. Although approximate, it efficiently captures attention-level epistemic uncertainty and serves as a practical surrogate for model-aware noise selection. $\dagger$$\dagger\daggerWe$ use “Bayesian” in the sense of epistemic uncertainty estimation, following BALD (Houlsby et al., 2011), rather than a full posterior derivation. As shown in Table 4, experimental validation confirms that this method is sufficient for selecting optimal noisy samples from the model’s perspective.
@@
5 return Generate video: $\hat{v}=\texttt{SampleVideo}(\mathbf{z}^{*},c,t)$;
@@ 3.3 Layer Selection via Cumulative BANSA Correlation
BANSA can be computed at any layer, but behavior differs across depth. Using all layers is costly, so we truncate at the smallest depth $d^{*}$ where the average BANSA up to $d$ remains highly correlated with the full-layer score. Given a noise seed $\mathbf{z}_{i}\in\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$ and $L$ attention layers, we compute per-layer scores ${\text{BANSA-E}}^{(l)}(\mathbf{z}_{i},c,t)$ and define the cumulative average up to layer $d$ as:
${\widehat{\text{BANSA-E}}}_{\leq d}(\mathbf{z}_{i},c,t):=\frac{1}{d}\sum_{l=1}^{d}{\text{BANSA-E}}^{(l)}(\mathbf{z}_{i},c,t).\vskip-5.0pt$ (7)
To find $d^{*}$, we compute the Pearson correlation (Pearson, 1895) between the partial average ${\widehat{\text{BANSA-E}}}_{\leq d}$ and the full-layer average ${\widehat{\text{BANSA-E}}}_{\leq L}$, and choose the smallest $d$ such that $\text{Corr}\!\left({\widehat{\text{BANSA-E}}}_{\leq d},\,{\widehat{\text{BANSA-E}}}_{\leq L}\right)\geq\tau.$ We validate this procedure using 100 prompts and 10 noise seeds across all evaluated models as detailed in Appendix D. As shown in Figure 7, the correlation quickly stabilizes at a moderate depth, allowing us to select $d^{*}$ without using all layers. We then define the BANSA score as ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ to guide noise selection, as summarized in Algorithm 1. With $d^{*}$ fixed per model, it adds no runtime cost, and ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ (truncated layer) consistently matches the full-layer score and robustly preserves generation quality across models, confirming its reliability as a practical surrogate as shown Appendix D and E (see detail).
4 Analysis of BANSA-Selected Noise
In this section, we aim to investigate the property of noise seeds selected by BANSA, and consist of three experiments which declare our framework and interpret the noise seeds. Our framework does not seek a universal “golden” seed, but selects the most suitable one for each prompt by minimizing uncertainty. The effectiveness of a noise seed is therefore prompt-dependent: a seed that improves one prompt may degrade another.
Experiment 1: Cross-Prompt Behavior.
Table 1: Cross-prompt evaluation of a high-BANSA seed from A.
From To Subject. BackGround